OpenStudy (anonymous):

A rectangle has a length of x+4 centimeters and a width of x centimeters. The perimeter of the rectangle is 28 centimeters. What is the value of x?

jigglypuff314 (jigglypuff314):

Perimeter = 2(length) + 2(width) so plug in (length) = x+4 and (width) = x and P = 28 so 28 = 2(x+4) + 2(x) do you know how to solve for x from there? :)

OpenStudy (anonymous):

no

jigglypuff314 (jigglypuff314):

well you would first distribute the 2 so 28 = 2(x+4) + 2(x) becomes 28 = 2x + 8 + 2x then add like terms (2x+2x) = 4x so 28 = 4x + 8 subtract 8 from both sides then divide both sides by 4

OpenStudy (anonymous):

5

jigglypuff314 (jigglypuff314):

correct :)